OpenStudy (anonymous):

If x^2+y^2=16 y^1 equals ?

OpenStudy (anonymous):

derivative, not solving for y

OpenStudy (mr.math):

You need to do implicit differentiation: \[2x+2yy'=16y' \implies y'=\frac{2x}{16-2y}=\frac{x}{8-y}.\]

OpenStudy (mertsj):

But the derivative of a constant is 0

OpenStudy (mertsj):

So the result should be -x/y

OpenStudy (anonymous):

@mr.math x^2+y^2=16 , derivative of y equals?

OpenStudy (mr.math):

Oh I thought it was \(x^2+y^2=16y\). If the problem is as you stated in your last comment, then: \(2x+2yy'=0 \implies y'=-\frac{x}{y}\), as Mertsj said.

OpenStudy (anonymous):

Thank you